OpenStudy (anonymous):

Solve this equation please h/6 =8

OpenStudy (anonymous):

Multiply both sides of the equation by 6.

OpenStudy (anonymous):

And m/-4=-5

OpenStudy (anonymous):

And m/-4=-5

OpenStudy (jamesj):

Multiply both sides by -4

OpenStudy (anonymous):

-4 x -4 and -5 x -4?

OpenStudy (anonymous):

m/-4 by -4, and -5 by -4.

OpenStudy (anonymous):

I don't get it.

OpenStudy (anonymous):

I don't get it.

OpenStudy (anonymous):

I just need to know what m =

OpenStudy (anonymous):

\[\frac{m}{-4} \times -4 = -5 \times -4\]

OpenStudy (anonymous):

I don't get it.

OpenStudy (anonymous):

but what does m equal?

OpenStudy (anonymous):

but what does m equal?

OpenStudy (anonymous):

\[\frac{m}{-4}\times -4= m \times \frac{-4}{-4} = m \times 1 = m\]\[m=-5 \times -4\]

OpenStudy (anonymous):

Thanks everyone my head is just not working for equations today.